OpenStudy (anonymous):

integral of dx over sqrtof e^(2x) - 1

OpenStudy (anonymous):

\[\int\limits_{\frac{ }{ }}^{} \frac{ dx }{ \sqrt{e{2x}-1} }\]

OpenStudy (anonymous):

Looks like you meant to type\[\int \dfrac{dx}{\sqrt{e^{2x}-1}}\] Let u =e^x du = e^x dx 1/e^x du = dx 1/u du = dx Using this substitution, the integral becomes \[\int \dfrac{\frac{1}{u} du}{\sqrt{u^2-1}}\\ \int \dfrac{du}{u\sqrt{u^2-1}}\] Look easier?

OpenStudy (anonymous):

\[\Large e^x=\sec\alpha \] Would also maybe help, because \[\Large \sec^2\alpha -1 = \tan^2 \alpha \]

OpenStudy (anonymous):

That substitution definitely works, too. In the end, it's all the same.

OpenStudy (anonymous):

\[\Large x= \ln (\sec\alpha) \\ \Large \frac{dx}{d\alpha}=\frac{1}{\sec\alpha}\cdot \tan\alpha \sec\alpha= \tan\alpha \]

OpenStudy (anonymous):

Was more a guess, I am actually a bit surprised myself, seems to cancel out very well (-:

OpenStudy (anonymous):

sitsh i did actually until there .. and another substitution is needed trigonometric substitution u=sint for example.. and i get \[\sqrt{-\cos^2 -1}\]

OpenStudy (anonymous):

The substitution you should have made was \[u=\sec t\\ du=\sec t \tan t \;dt\] The sub you used should be used if you're given an integral containing \[\sqrt{1-u^2}.\]

OpenStudy (anonymous):

i em tryin to find that but at the end i get \[\sqrt{-\cos^2t}\]

OpenStudy (anonymous):

and it doestn work so more

OpenStudy (anonymous):

thanks .. i did.. with u=sect..